# Beyond English: Evaluating Automated Measurement of Moral Foundations in Non-English Discourse with a Chinese Case Study

### v2

@@ Abstract
This study explores computational approaches for measuring moral foundations (MFs) in non-English corpora. Since most resources are developed primarily for English, cross-linguistic applications of moral foundation theory remain limited. Using Chinese as a case study, this paper evaluates the effectiveness of applying English resources to machine translated text, local language lexicons, multilingual language models, and large language models (LLMs) in measuring MFs in non-English texts. The results indicate that machine translation and local lexicon approaches are insufficient for complex moral assessments, frequently resulting in a substantial loss of cultural information. In contrast, multilingual models and LLMs demonstrate reliable cross-language performance with transfer learning, with LLMs excelling in terms of data efficiency. Importantly, this study also underscores the need for human-in-the-loop validation of automated MF assessment, as the most advanced models may overlook cultural nuances in cross-language measurements. The findings highlight the potential of LLMs for cross-language MF measurements and other complex multilingual deductive coding tasks.
@@ Introduction
Moral intuitions have long fascinated social scientists, as they help explain a wide range of cognitive and behavioral phenomena across individuals and groups (Effron and Helgason, 2022). Moral foundation theory (MFT) is among the most prominent psychology frameworks for understanding the origin and development of human morality (Graham et al., 2013). Rooted in moral nativism, MFT argues there are five universal moral foundations—care/harm, fairness/cheating, authority/subversion, loyalty/betrayal, and sanctity/degradation—that transcend languages and cultures and underlie people’s moral judgments and decision-making processes (Graham et al., 2013)  [^1] Each MF includes both virtue and vice dimensions. This paper uses the virtue label to represent the entire foundation.. While some scholars propose other foundations (Haidt, 2012; Atari et al., 2023), these five foundations have received the most empirical validation across domains, languages, and cultures (Iurino and Saucier, 2020).
@@
A growing body of literature employs MFT to investigate online social behaviors. For example, MFT provides a framework for understanding rising political polarization. Individuals who prioritize loyalty, authority, and sanctity foundations are more likely to endorse conservative views and engage in polarized political discourse, while those affiliated with liberal ideologies tend to value all MFs more evenly (Koleva et al., 2012; Haidt and Graham, 2007). MFT also sheds light on online cultural clashes. Atran (2007) identified differing valuations of the sanctity foundation as a key factor in many religious and ideological conflicts. Beyond polarization, MFT has been applied to a range of social issues, including climate change (Markowitz and Shariff, 2012), vaccine hesitancy (Amin et al., 2017), anti-abortion views (Koleva et al., 2012), nationalism (Kertzer et al., 2014), violent collectivism (Nussio, 2023), and terrorism (Tamborini et al., 2020).
@@
Given its broad relevance, measuring MF values in online discourse is essential; yet, automated extraction of MF values from large-scale texts remains challenging, particularly for non-English corpora. Like other latent human values, MFs are often conveyed through abstract narratives that vary across languages. Also, most computational resources for the measurement of MFs are designed for English content (Hoover et al., 2020; Trager et al., 2022). This reliance on English hinders cross-cultural comparative research on MFT and limits theoretical advancement from non-English data (Cheng and Zhang, 2023). Although MFT is a universal theory intended to apply across languages and cultures, the limited availability of non-English resources severely restricts its research scope and further development of the theoretical framework (Graham et al., 2013).
@@
In this work, we investigate various computational approaches for cross-language measurement of MFs with a particular focus on data-efficiency. We use Chinese as an example and find that (1) MF local language lexicons yield suboptimal performance. They are worse than machine translation approaches that utilize established English measurements such as Mformer; (2) multilingual language models can achieve moderate success when trained on annotated English data along with some local language labeled data. However, this strategy is less data-efficient in measuring MFs than in other deductive coding tasks (e.g., Röttger et al., 2022). (3) LLMs outperform other approaches in cross-language MF measurements in both accuracy and data efficiency. Simply fine-tuning and augmenting with English annotated data can achieve strong performance in non-English corpora; (4) nevertheless, this performance is in consistent on different MF values, as LLMs may overlook cultural nuances in cross-language measurements, particularly for culturally distinct values.  [^2] Code is available at GitHub https://github.com/calvinchengyx/cross-lan-mft-measure.
@@
The unique link between word usage and the expressed moral values provides a theoretical foundation for automated MF measurement from online texts (Brady, Crockett, and Van Bavel, 2020; Gantman and Van Bavel, 2014, 2016). Scholars have explored different computational approaches, including dictionaries (Graham, Haidt, and Nosek, 2009; Hopp et al., 2021), word embeddings (Kwak et al., 2021; Araque, Gatti, and Kalimeri, 2020), machine learning (Lan and Paraboni, 2022), deep learning language models (Preniqi et al., 2024; Nguyen et al., 2024) and LLMs (Rathje et al., 2024). These computational methods demonstrate great advantages on scalability and labor intensity compared to traditional human annotations.
However, these approaches are primarily developed for English language, and are not directly applicable to non-English texts due to several major concerns: differences in cultural contexts, the lack of annotated datasets, and limitations in domain generalizability. To address the cross-language challenges and bridge knowledge gaps in MF measurements, various computational approaches have been proposed, which can be broadly categorized into two paths: machine translation to English and the development of cross-language measurement tools (Zhuang et al., 2020).
@@
Translation is a widely used technique in cross-language MF measurement. For instance, the MFT survey has been translated into over 20 languages for cross-lingual studies (Yilmaz et al., 2016; Nilsson and Erlandsson, 2015). For large scale text analysis, the development of multilingual neural machine translation has significantly improved translation quality compared to earlier statistical methods, enhancing context understanding, ambiguity resolution, and fluency (Stasimioti et al., 2020). This advancement enables a machine-translated approach to cross-language MF measurement by translating target languages into English and applying established English-based methods (Artetxe, Labaka, and Agirre, 2020).
Previous research has proposed various established methods for automatically measuring MFs in English, including dictionaries (Graham, Haidt, and Nosek, 2009; Frimer et al., 2019; Hopp et al., 2021), word embeddings (Kwak et al., 2021; Araque, Gatti, and Kalimeri, 2020), machine learning and deep learning models (Preniqi et al., 2024; Nguyen et al., 2024) trained on annotated English-language social media data (Hoover et al., 2020; Trager et al., 2022).
@@
Word-count methods with crafted English moral lexicons are common. Scholars have developed four English MFDs. The original MFD is an expert-crafted dictionary containing a list of 600 words across five foundation values (Graham, Haidt, and Nosek, 2009). Frimer et al. (2019) then expanded this vocabulary to MFD2 with over 2,000 words by automatically querying similar words with word2vec word embeddings (Mikolov et al., 2013). Similarly, Araque, Gatti, and Kalimeri (2020) extended the original MFD to a MoralStrength dictionary with approximately 1,000 English lemmas based on WordNet synsets (Princeton, 2010). Compared to MFD2, MoralStrength added a round of crowd-sourced ratings on expanded lemmas. Hopp et al. (2021), however, curated a fully crowd-sourced dictionary named eMFD. It is different from previous expert-curated dictionaries for its layperson focus, contextual annotations, probability labeling and large vocabulary with 3,200 words.
@@
Semantic similarity methods using embeddings are another approach. To address the limitations of word-count methods, such as context insensitivity and vocabulary coverage (Nguyen et al., 2024), scholars have introduced semantic similarity methods. For example, Kwak et al. (2021) proposed an embedding framework FrameAxis. It predefines a vector space of micro moral frames with two sets of opposing seed words. Target documents are then converted into vectors using word embedding models, and their MF values are determined by comparing to the micro-frames. This method has been used to extract MF values from various online texts (e.g., Mokhberian et al., 2020; Jing and Ahn, 2021).
@@
Supervised classification models with annotated English-language training data have also been used. With advancements in language models and efforts to create human-labeled MF training datasets (Hoover et al., 2020; Trager et al., 2022), recent studies have demonstrated the potential of fine-tuning language models. For example, Preniqi et al. (2024) fine-tuned a BERT-based classifier MoralBert with large-scale annotated English data and achieved state-of-the-art performance. To address generalizability limitations in out-of-domain datasets (Liscio et al., 2022), Nguyen et al. (2024) proposed another language model Mformer, and reported superior performance compared to other established English methods in evaluations.
Although machine translation offers several advantages in cross-language MF measurement, including interpretability, scalability, efficiency and accessibility, it also faces significant limitations. First, translation quality varies across languages and domains (Ranathunga et al., 2023). In some low-resource languages like Tamil, machine translation often makes errors in translating domain terms, polysemous words, and contains repetitions for semantically similar terms (Ramesh et al., 2021). Second, it often fails to retain non-propositional information, such as emotion nuances. This can lead to emotion loss, toning down, or amplification across languages, introducing bias in subsequent analyses (Troiano, Klinger, and Padó, 2020). Third, machine translation struggles to capture cultural elements, which is a major concern for cross-cultural and comparative research (Haidt, 2012). It often shows limited performance with rare or culture-specific words, idiomatic phrases, and metaphor recognition (Dorothy, Keiko, and Vladimir, 2019). Thus, it remains unclear whether machine translation is a reliable method for measuring cross-language MFs.
@@
A second path is to develop cross-language tools, where scholars create computational MF resources tailored to local languages. Common cross-language measurements include local language dictionaries, task-specific multilingual language models, and LLMs.
@@
Due to the efficiency at scale and multilingual capabilities, local language dictionaries are widely used to estimate MF values from non-English texts (Hopp et al., 2021). Developing local language MFDs generally involves three steps: (1) translating English MFDs to target languages; (2) adding culturally specific and non-translatable vocabulary; and (3) validating with native speakers and local language corpora. Several extensive non-English MFDs have been developed and validated in Turkish (Alper et al., 2020), Japanese (Matsuo et al., 2019), Portuguese (Carvalho et al., 2020), and Chinese (Cheng and Zhang, 2023). Despite the abovementioned advantages, the dictionary approach still faces the inherent limitations of general bag-of-words methods. In this paper, we use C-MFD2—a Chinese MFD—as a cross-language tool to evaluate a local language dictionary approach. We also test a semantic similarity approach using the FrameAxis architecture and cross-language word embedding models.
Multilingual Language Models
To overcome the limitations of bag-of-words methods, literature has suggested machine learning and deep learning approaches. A primary challenge with these methods is the scarcity of annotated data in local languages for model training (e.g., Ji et al., 2024; Nguyen et al., 2024). Therefore, scholars have adopted transfer learning techniques that leverage English-annotated resources for cross-language classifier development. Two major transfer learning strategies are commonly proposed: (1) machine-translating annotated English-language data into local languages to train monolingual models (Schuster et al., 2019), or (2) using annotated English-language data to train multilingual models (Barriere and Balahur, 2020). Given the strong performance of multilingual language models in deductive coding tasks such as sentiment analysis (e.g., Barriere and Balahur, 2020) and hate speech detection (e.g., Röttger et al., 2022), this paper focuses on the second transfer learning strategy and evaluates multilingual language models for cross-language MF measurement.
@@
The rise of LLMs provides an alternative for cross-language MF measurement. LLMs show exceptional zero/few-shot learning capability, enabling them to directly label human values out-of-the-box, which is particularly valuable for tasks with limited human annotated data (Ziems et al., 2024). They also have demonstrated strong performance in measuring various human values, including emotions, stance, and political ideology (Ziems et al., 2024; Gilardi, Alizadeh, and Kubli, 2023; Törnberg, 2023). Notably, LLMs sometimes are not as good as specialized fine-tuned language models (Amin, Cambria, and Schuller, 2023; Preniqi et al., 2024), which may be due to the lack of explicit, colloquial definitions of the target human values (Ziems et al., 2024). MFT’s well-established conceptual framework may help address this limitation. Not only are LLMs pre-trained on rich MFT literature (e.g., Abdulhai et al., 2023), but MFT also offers clear guidance for crafting clear and effective prompts. Additionally, LLMs trained on vast multilingual data exhibit promising capabilities to handle cross-language measurements (Ahuja et al., 2023).
Despite the strengths in accessibility, efficiency, multilingualism, and reasoning, there are also some concerns for LLMs’ performance in cross-language MF measurement. First, LLMs exhibit a substantial degree of subordinate multilingualism, displaying proficiency in some languages but not others (Zhang et al., 2023), which has a strong correlation with the proportion of those languages in the pre-training corpus (Li et al., 2024). Second, there are potential language biases, particularly in human-value relevant coding tasks (Kirk et al., 2024). For example, non-English prompts are more likely to generate malicious responses compared to English prompts (Shen et al., 2024). LLMs may also have moral bias across languages, as pre-trained multilingual language models often display distinct moral directions across languages (Hämmerl et al., 2023). Third, different LLMs exhibit varying baseline moral tendencies (Ji et al., 2024). For instance, GPT-3’s MF preferences align more closely with politically conservative individuals when minimal prompt engineering is used in zero-shot learning (Abdulhai et al., 2023). These moral tendencies, however, are highly sensitive to prompting, with different prompting strategies significantly influencing classification outcomes in MF measurements (Abdulhai et al., 2023).
Thus, it is unclear how LLMs performs in cross-language MF measurement tasks. This paper selects a cutting-edge, open-source model—Llama3.1 (Meta, 2024) to evaluate the LLMs approach.
@@
Table 1 shows details of the benchmarking and training datasets used in this work. We used three human annotated datasets—moral foundation vignettes (MFV), Chinese moral scenarios (CCS) and Chinese core values (CCV), to benchmark the performance of different cross-language MF measurement approaches. And we used three English annotated MFT datasets—Reddit Corpus, Twitter Corpus and news corpus—for model training and fine-tuning, aligning with the work of Nguyen et al. (2024).
@@
MFV is a list of social behaviors constructed by psychologists based on MFT, describing the violations of specific moral values from a third-party perspective (Clifford et al., 2015). It has been widely validated and used to assess measures of moral judgment (e.g., Atari et al., 2023; Curry, Chesters, and Van Lissa, 2019). The original MFV is in English; we manually translate it to Chinese and ensure the cultural context is preserved. MFV is used as a universal stimulus to benchmark the performance of different cross-language MF measurement approaches on an expert-crafted sample.
@@
CCS is list of moral scenarios written by laypeople to describe their intuitive understanding of MF values (Cheng and Zhang, 2023). It is written by 202 native Chinese speakers describing situations that either violate or respect specific MF values, with minimal training in MFT literacy. This reverse-annotation method is commonly used for validating MF measurements (e.g., Cheng and Zhang, 2023; Frimer et al., 2019; Matsuo et al., 2019). It incorporates culturally specific content, reflecting native speakers’ natural and intuitive understanding of MFT in a real-word context.
@@
CCV is a human-annotated real-world dataset, including 6,994 sentences collected from four local Chinese news websites. [^3] CCV includes data from the CMOS corpus (Peng et al., 2021), China Cultural and Ethical Website wenming.cn, Youth Patriotism News agzy.youth.cn, and Sohu News news.sohu.com. The dataset is annotated by three native Chinese speakers based on the Chinese core socialist moral value coding scheme, which is highly correlate with the five universal moral foundation values (Liu et al., 2022). The original CCV dataset includes eight labels, [^4] The Chinese core socialist moral values include civility, justice, equality, rule of law, patriotism, dedication, integrity, and friendship. The curated mapping scheme to the five moral foundation values is in the Appendix which are qualitatively evaluated and re-mapped to the universal five foundation values following an expert-curated mapping scheme. We sample 20% of the CCV as the primary benchmarking dataset to evaluate the performance across approaches stratifying on the values. The remaining 80% is used as local-language annotated data to test the data-efficiency fine-tuning language models.
@@
We use three English annotated MF datasets for transfer learning, including a Twitter corpus (Hoover et al., 2020), a Reddit corpus (Trager et al., 2022) and a news corpus (Hopp et al., 2021), which are widely used in training English MF classifiers and show reliable performance (Nguyen et al., 2024; Preniqi et al., 2024).

### v3

@@ Abstract
This study explores computational approaches for measuring moral foundations (MFs) in non-English corpora. Since most resources are developed primarily for English, cross-linguistic applications of moral foundation theory remain limited. Using Chinese as a case study, this paper evaluates the effectiveness of applying English resources to machine translated text, local language lexicons, multilingual encoder-only language models, and decoder-only large language models (LLMs) in measuring MFs in non-English texts. The results indicate that machine translation and local lexicon approaches are insufficient for complex moral assessments, frequently resulting in a substantial loss of cultural information. In contrast, language models demonstrate reliable cross-language performance with transfer learning, with LLMs excelling in terms of data efficiency. Importantly, this study also underscores the need for human-in-the-loop validation of automated MF assessment, as even the most advanced models may overlook cultural nuances and face potential risks in cultural misalignment. The findings highlight the potential of LLMs for cross-language MF measurements and other complex multilingual deductive coding tasks.
@@ Introduction
Moral intuitions have long fascinated social scientists, as they help explain a wide range of cognitive and behavioral phenomena across individuals and groups (Effron and Helgason 2022). Moral foundation theory (MFT) is among the most prominent psychology frameworks for understanding the origin and development of human morality (Graham et al. 2013). Rooted in moral nativism, MFT argues there are five universal moral foundations—care/harm, fairness/cheating, authority/subversion, loyalty/betrayal, and sanctity/degradation—that transcend languages and cultures and underlie people’s moral judgments and decision-making processes (Graham et al. 2013). [^1] Each MF includes both virtue and vice dimensions. We use the virtue label to represent the foundation value. While some scholars propose other foundations (Haidt 2012; Atari et al. 2023a), these five foundations have received the most empirical validation across domains, languages, and cultures (Iurino and Saucier 2020).
@@
A growing body of literature employs MFT to investigate online social behaviors. For example, MFT provides a framework for understanding rising political polarization. Individuals who prioritize loyalty, authority, and sanctity foundations are more likely to endorse conservative views and engage in polarized political discourse, while those affiliated with liberal ideologies tend to value all MFs more evenly (Koleva et al. 2012; Haidt and Graham 2007). MFT also sheds light on online cultural clashes. Atran (2007) identified differing valuations of the sanctity foundation as a key factor in many religious and ideological conflicts. Beyond polarization, MFT has been applied to a range of social issues, including climate change (Markowitz and Shariff 2012), vaccine hesitancy (Amin et al. 2017), anti-abortion views (Koleva et al. 2012), nationalism (Kertzer et al. 2014), collective violence (Nussio 2023), and terrorism (Tamborini et al. 2020).
@@
Given its broad relevance, measuring MF values in online discourse is essential; yet, automated extraction of MF values from large-scale texts remains challenging, particularly for non-English corpora. Like other latent human values, MFs are often conveyed through abstract narratives that vary across languages. Also, most computational resources for the measurement of MFs are designed for English content (Hoover et al. 2020; Trager et al. 2022). This reliance on English hinders cross-cultural comparative research on MFT and limits theoretical advancement from non-English data (Cheng and Zhang 2023). Although MFT is intended to apply across languages and cultures, the limited availability of non-English resources severely restricts its research scope and further development of the theoretical framework (Graham et al. 2013).
@@
In this work, we investigate various computational approaches for cross-language measurement of MFs with a particular focus on data-efficiency. We use Chinese as an example and find that (1) MF local language lexicons yield suboptimal performance. They are worse than machine translation approaches that utilize established English measurements such as Mformer. (2) Multilingual encoder-only models can achieve moderate success when trained on annotated English data along with some local language labeled data. However, this strategy is less data-efficient in measuring MFs than in other deductive coding tasks, such as hate speech detection (Röttger et al. 2022). (3) Decoder-only LLMs outperform other approaches in cross-language MF measurements in both accuracy and data efficiency. Simply fine-tuning and augmenting with English annotated data can achieve strong performance on non-English corpora; (4) nevertheless, this performance is inconsistent on different MF values, as LLMs may overlook cultural nuances in cross-language measurements, particularly for culturally distinct values  [^2] Code is available on the project’s GitHub https://github.com/calvinchengyx/cross-lan-mft-measure.
@@
The unique link between word usage and the expressed moral values provides a theoretical foundation for automated MF measurement from online texts (Brady, Crockett, and Van Bavel 2020; Gantman and Van Bavel 2014, 2016). Scholars have explored different computational approaches, including dictionaries (Graham, Haidt, and Nosek 2009; Hopp et al. 2021), word embeddings (Kwak et al. 2021; Araque, Gatti, and Kalimeri 2020), machine learning (Lan and Paraboni 2022), deep learning language models (Preniqi et al. 2024; Nguyen et al. 2024) and LLMs (Rathje et al. 2024). These computational methods demonstrate great advantages on scalability and labor intensity compared to traditional human annotations.
However, these approaches are primarily developed for English, and are not directly applicable to non-English texts due to several major concerns: differences in cultural contexts, the lack of annotated datasets, and limitations in domain generalizability. To address the cross-language challenges and bridge knowledge gaps in MF measurements, various computational approaches have been proposed, which can be broadly categorized into two paths: machine translation to English and the development of cross-language measurement tools (Zhuang et al. 2020).
@@
Translation is a widely used technique in cross-language MF measurement. For instance, the MFT survey has been translated into over 20 languages for cross-lingual studies (Yilmaz et al. 2016; Nilsson and Erlandsson 2015). For large scale text analysis, the development of multilingual neural machine translation has significantly improved translation quality compared to earlier statistical methods, enhancing context understanding, ambiguity resolution, and fluency (Stasimioti et al. 2020). This advancement enables a machine-translated approach to cross-language MF measurement by translating target languages into English and applying established English-based methods (Artetxe, Labaka, and Agirre 2020).
Established methods for automatically measuring MFs in English include dictionaries (Graham, Haidt, and Nosek 2009; Frimer et al. 2019; Hopp et al. 2021), word embeddings (Kwak et al. 2021; Araque, Gatti, and Kalimeri 2020), machine learning and deep learning models (Preniqi et al. 2024; Nguyen et al. 2024) trained on annotated English-language social media data (Hoover et al. 2020; Trager et al. 2022).
@@
Word-count methods with crafted English moral lexicons are common. There are four common English MFDs. The original MFD is an expert-crafted dictionary containing a list of 600 words across five foundation values (Graham, Haidt, and Nosek 2009). Frimer et al. (2019) then expanded this vocabulary to MFD2 to over 2,000 words by automatically identifying similar words with word2vec word embeddings (Mikolov et al. 2013). Similarly, Araque, Gatti, and Kalimeri (2020) extended the original MFD to a MoralStrength dictionary with approximately 1,000 English lemmas based on WordNet synsets (Princeton 2010). Compared to MFD2, MoralStrength added a round of crowd-sourced ratings on the expanded lemmas. Hopp et al. (2021), however, curated a fully crowd-sourced dictionary named eMFD. It is different from previous expert-curated dictionaries for its layperson focus, contextual annotations, probability labeling and large vocabulary with 3,200 words.
@@
Semantic similarity methods using embeddings are another approach. To address the limitations of word-count methods, such as context insensitivity and vocabulary coverage (Nguyen et al. 2024), scholars have introduced semantic similarity methods. For example, Kwak et al. (2021) proposed an embedding framework FrameAxis. It predefines a vector space of micro moral frames with two sets of opposing seed words. Target documents are then converted into vectors using word embedding models, and their MF values are determined by comparing to the micro-frames. This method has been used to extract MF values from various online texts (e.g., Mokhberian et al. 2020; Jing and Ahn 2021).
@@
Supervised classification models with annotated English-language training data have also been used. With advancements in language models and efforts to create human-labeled MF training datasets (Hoover et al. 2020; Trager et al. 2022), recent studies have demonstrated the potential of fine-tuning language models. For example, Preniqi et al. (2024) fine-tuned a BERT-based classifier MoralBert with large-scale annotated English data and achieved state-of-the-art performance. To address generalizability limitations in out-of-domain datasets (Liscio et al. 2022), Nguyen et al. (2024) proposed another language model Mformer, and reported superior performance compared to other established English methods in evaluations.
Although machine translation offers several advantages in cross-language MF measurement, including interpretability, scalability, efficiency and accessibility, it also faces significant limitations. First, translation quality varies across languages and domains (Ranathunga et al. 2023). In some low-resource languages like Tamil, machine translation often makes errors in translating domain terms, polysemous words, and contains repetitions for semantically similar terms (Ramesh et al. 2021). Second, it often fails to retain non-propositional information, such as emotional nuances. This can lead to a loss of emotions, toning down, or amplification across languages, introducing bias in subsequent analyses (Troiano, Klinger, and Padó 2020). Third, machine translation struggles to capture cultural elements, which is a major concern for cross-cultural and comparative research (Haidt 2012). It often shows limited performance with rare or culture-specific words, idiomatic phrases, and metaphor recognition (Dorothy, Keiko, and Vladimir 2019). Thus, it remains unclear whether machine translation is a reliable method for measuring cross-language MFs.
@@
A second path is to develop cross-language tools, where scholars create computational MF resources tailored to local languages. Common cross-language measurements include local language dictionaries, task-specific encoder-only language models, and LLMs.
@@
Due to the efficiency at scale and multilingual capabilities, local language dictionaries are widely used to estimate MF values from non-English texts (Hopp et al. 2021). Developing local language dictionaries generally involves three steps: (1) translating English dictionaries to target languages; (2) adding culturally specific and non-translatable vocabulary; and (3) validating with native speakers and local language corpora. Several extensive non-English MFDs have been developed and validated in Turkish (Alper et al. 2020), Japanese (Matsuo et al. 2019), Portuguese (Carvalho et al. 2020), and Chinese (Cheng and Zhang 2023). Despite the abovementioned advantages, the dictionary approach still faces the inherent limitations of general bag-of-words methods. In this paper, we use C-MFD2—a Chinese MFD—as a cross-language tool to evaluate a local language dictionary approach. We also test a semantic similarity approach using the FrameAxis architecture and cross-language word embedding models.
Multilingual encoder-only models
To overcome the limitations of bag-of-words methods, literature has suggested machine learning and deep learning approaches. A primary challenge with these methods is the scarcity of annotated data in local languages for model training (e.g., Ji et al. 2024; Nguyen et al. 2024). Therefore, scholars have adopted transfer learning techniques that leverage English-annotated resources for cross-language classifier development. Two major transfer learning strategies are commonly proposed: (1) machine-translating annotated English-language data into local languages to train monolingual models (Schuster et al. 2019), or (2) using annotated English-language data to train multilingual encoder-only models (Barriere and Balahur 2020). Multilingual models have demonstrated strong performance in deductive coding tasks, such as sentiment analysis (e.g., Barriere and Balahur 2020) and hate speech detection (e.g., Röttger et al. 2022). Nevertheless, they also have shown language bias in morality classification tasks, as pre-trained multilingual encoder-only language models often display distinct moral directions across languages (Hämmerl et al. 2023). This paper focuses on the second transfer learning strategy and evaluates multilingual encoder-only models for cross-language MF measurement.
@@
The rise of decoder-only LLMs provides an alternative for cross-language MF measurement. LLMs show exceptional zero/few-shot learning capability, enabling them to directly label human values out-of-the-box, which is particularly valuable for tasks with limited human annotated data (Ziems et al. 2024). They also have demonstrated strong performance in measuring various human values, including moral reasoning tasks (Ziems et al. 2024; Agarwal et al. 2024). Notably, LLMs sometimes are not as good as specialized fine-tuned language models (Amin, Cambria, and Schuller 2023; Preniqi et al. 2024), which may be due to the lack of explicit, colloquial definitions of the target human values (Ziems et al. 2024). MFT’s well-established conceptual framework may help address this limitation. Not only are LLMs pre-trained on rich MFT literature (e.g., Abdulhai et al. 2023), but MFT also offers clear guidance for crafting clear and effective prompts. Additionally, LLMs trained on vast multilingual data exhibit promising capabilities to handle cross-language measurements (Ahuja et al. 2023).
Despite the strengths in accessibility, efficiency, multilingualism, and reasoning, there are also some concerns in using LLMs’ in cross-language MF measurement. First, LLMs exhibit a substantial degree of subordinate multilingualism, displaying proficiency in some languages but not others (Zhang et al. 2023), which has a strong correlation with the proportion of those languages in the pre-training corpus (Li et al. 2024). Second, there are potential language biases, particularly in human-value relevant coding tasks (Kirk et al. 2024; Yu et al. 2024). For example, non-English prompts are more likely to generate malicious responses compared to English prompts (Shen et al. 2024). Third, different LLMs exhibit varying baseline moral tendencies (Ji et al. 2024). For instance, GPT-3’s MF preferences align more closely with politically conservative individuals when minimal prompt engineering is used in zero-shot learning (Abdulhai et al. 2023). These moral tendencies, however, are sensitive to prompting, with different prompting strategies significantly influencing classification outcomes in MF measurements (Abdulhai et al. 2023; Agarwal et al. 2024).
Thus, it is unclear how LLMs performs in cross-language MF measurement tasks. This paper selects a cutting-edge, open-source model—Llama3.1 (Meta 2024) to evaluate on LLM approach.
@@
Table 1 shows details of the benchmarking and training datasets used in this work. We used three human annotated datasets—moral foundation vignettes (MFV), Chinese moral scenarios (CCS) and Chinese core values (CCV), to benchmark the performance of different cross-language MF measurement approaches. We also used three English annotated MFT datasets for model training and fine-tuning.
@@
MFV is a list of social behaviors constructed by psychologists based on MFT, describing the violations of specific moral values from a third-party perspective (Clifford et al. 2015). It has been widely validated and used to assess measures of moral judgment (e.g., Graham, Haidt, and Nosek 2009; Kivikangas et al. 2021; Ji et al. 2024). We used the MFV as an expert-crafted benchmark stimulus to evaluate the baseline performance of different moral foundation measurement approaches. Since the original MFV is in English, we followed a careful translation process to ensure cultural neutrality in Chinese. First, a native Chinese speaker translated the vignettes with minor modifications to preserve cultural appropriateness. Then, two additional native speakers reviewed the translations to assess whether the scenarios remained representative and meaningful in Chinese cultural contexts. This ensured that the translated vignettes could serve as a culturally neutral benchmark for cross-method comparability. [^3] The translated vignettes are available on the project’s GitHub page.
@@
CCS is list of moral scenarios written by 202 native Chinese to describe their intuitive understandings of MF values (Cheng and Zhang 2023). This reverse-annotation method is commonly used for validating MF measurements (e.g., Cheng and Zhang 2023; Frimer et al. 2019; Matsuo et al. 2019). It incorporates culturally specific content, reflecting native speakers’ natural and intuitive understanding of MFT in a real-word context.
@@
CCV is a human-annotated real-world dataset, including 6,994 sentences collected from four local news websites. [^4] CCV includes data from the CMOS corpus (Peng et al. 2021), China Cultural and Ethical Website wenming.cn, Youth Patriotism News agzy.youth.cn, and Sohu News news.sohu.com The dataset is annotated by three native Chinese speakers based on the Chinese core socialist moral value coding scheme, which is highly correlate with the five universal moral foundation values (Liu et al. 2022). The original CCV dataset includes eight labels, [^5] The Chinese core socialist moral values include civility, justice, equality, rule of law, patriotism, dedication, integrity, and friendship. Table 3 in the Appendix showed the curated mapping scheme. we employed five Chinese native speakers with postgraduate degrees to re-label the action categories in CCV to five moral foundation values, excluding vice and virtue. The mapped values were determined by majority vote. We sample 20% of the CCV as the primary benchmarking dataset to evaluate the performance across approaches stratifying on the values. The remainder is used as training data to test the data-efficiency fine-tuning language models.
@@
We use three English annotated MF datasets for transfer learning, including a Twitter corpus (Hoover et al. 2020), a Reddit corpus (Trager et al. 2022) and a news corpus (Hopp et al. 2021), which are widely used in training English MF classifiers and show reliable performance (Nguyen et al. 2024; Preniqi et al. 2024).